OpenStudy (anonymous):

A jar is filled with 1,000 jellybeans. Each day, you eat 3 jellybeans. If you graphed the number of jellybeans in the jar as the output, compared to the day number as input, which of the following would not be true about the graph?

OpenStudy (anonymous):

A. The graph would be a line. B. The graph would rise from left to right. C. The graph would be only in the first quadrant. D. The larger the number of days, the fewer jellybeans there would be.

OpenStudy (tanner23456):

y = 1000 - 3x There fore the graph is linear so A is correct. The graph falls from left to right so B is not correct. The graph is only in the first quadrant because you can't have negative jelly beans. C is correct. More days = less jelly beans, D is correct. Correct: A, C, D Incorrect: B 

OpenStudy (anonymous):

oooooh ok

OpenStudy (tanner23456):

use y= mx+b as a guide to set up problems like this. y= 1000-3x could also be written as y=-3x+1000